A random variable T has a probability density function given by

f(t)={1/2,0<t<1 both inclusive

{x-1/2,1<t<c both inclusive

{0,elsewhere

where c is a positive constant

  1. sketch f(t) hence using geometric shapes or otherwise show that

c=(1+ 51/2)/2

2.Write down the median of T and the mode of T hence describe the skewness of the distribution(Give a reason for your answer)


1
Expert's answer
2021-07-26T06:44:43-0400

1.


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=1"




"\displaystyle\int_{0}^{1}\dfrac{1}{2}dx+\displaystyle\int_{1}^{c}(x-\dfrac{1}{2})dx=\dfrac{1}{2}[x]\begin{matrix}
 1 \\
 0
\end{matrix}+\dfrac{1}{2}[x^2-x]\begin{matrix}
 c \\
 1
\end{matrix}"

"=\dfrac{1}{2}+\dfrac{1}{2}(c^2-c-(1-1))=1"

"c^2-c-1=0"


"c=\dfrac{1\pm\sqrt{5}}{2}"

Since "c\geq 1," we take "c=\dfrac{1+\sqrt{5}}{2}."

"f(x) = \begin{cases}
 \dfrac{1}{2}, &0\leq x\leq 1 \\
 x-\dfrac{1}{2}, &1\leq x\leq\dfrac{1+\sqrt{5}}{2}\\
0, & elsewhere
\end{cases}"



2. The mode of a continuous probability distribution is the point at which the probability density function attains its maximum value.


"mode=\dfrac{1+\sqrt{5}}{2}>1"




"F(x) = \begin{cases}
 0, & x<0 \\
 \dfrac{1}{2}x, & 0\leq x<1 \\
 \dfrac{x^2}{2}-\dfrac{1}{2}x+ \dfrac{1}{2}, &1\leq x<\dfrac{1+\sqrt{5}}{2}\\
1, & x\geq \dfrac{1+\sqrt{5}}{2}\\
\end{cases}"

The median of a continuous probability distribution


"F(median)= \dfrac{1}{2}"

"median= 1"

"mean=\displaystyle\int_{-\infin}^{\infin}xf(x)dx="

"=\displaystyle\int_{0}^{1}\dfrac{1}{2}xdx+\displaystyle\int_{1}^{(1+\sqrt{5})/2}x(x-\dfrac{1}{2})dx"

"=\dfrac{1}{2}[\dfrac{1}{2}x^2]\begin{matrix}
 1 \\
 0
\end{matrix}+[\dfrac{1}{3}x^3-\dfrac{1}{4}x^2]\begin{matrix}
 (1+\sqrt{5}/2 \\
 1
\end{matrix}"

"=\dfrac{1}{4}+\dfrac{1}{2}(\dfrac{1+\sqrt{5}}{2})^3-\dfrac{1}{2}(\dfrac{1+\sqrt{5}}{2})^2-\dfrac{1}{3}+\dfrac{1}{4}"

"=\dfrac{11+5\sqrt{5}}{24}\approx0.92<1"

We have a negative (left) skewness


"mean<median<mode"